# Nonreciprocal Transmission in Hybrid Atomic Ensemble-Optomechanical Systems

### v1

@@ III Results and Discussion
In this section, we analyze the conditions under which optical nonreciprocity can be achieved in the hybrid optomechanical system. We investigate the effects of key system parameters such as coupling strengths, detuning, dissipation, and phase differences on the transmission properties. The results reveal how tuning these parameters allows for control over nonreciprocal transmission, offering potential applications in optical isolation and signal routing.
@@ III.1 Optical Nonreciprocal Transmission
The transmission amplitudes $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ are defined to delineate the optical signal transmission from cavity 1 to cavity 2 and vice versa. These transmission coefficients are derived from the input-output relations and exhibit a strong dependence on the system parameters, particularly the phase differences between the effective optomechanical couplings and the coupling strengths of the atomic ensemble. Figure 2 displays the dependence of transmission amplitudes $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ on the phase difference $\theta$ between the effective optomechanical couplings and the phase $\phi$ associated with the atomic ensemble. Notably, although the transmission amplitudes are not markedly influenced by phase $\phi$, both phases play a crucial role in the emergence of nonreciprocal transmission.
Figure 2: Density plot of transmission amplitudes ($\mathcal{T}_{12}$ and $\mathcal{T}_{21}$) versus relative phase difference $\theta$ between effective optomechanical coupling strength and the phase $\phi$. We have used $G_{1}/\gamma=0.5$, $G_{2}=G_{1}$, $f/\gamma=10$, $\kappa_{1}/\gamma=1$, $\kappa_{2}=\kappa_{1}$, $J_{1}/\gamma=0.5$, $\omega_{m}/\gamma=10$, $\Delta^{\prime}_{1}=\Delta^{\prime}_{2}=\Delta_{en}=\omega_{m}$, $|J_{2}|/\gamma=0.01$, and $J_{3}/\gamma=4.476e^{i\frac{\pi}{2}}$.
The results in Figure 2 suggest that optimal nonreciprocal transmission is realized when the coupling constant between the atomic ensemble and the mechanical oscillator, $J_{3}$, is complex. The complex nature of $J_{3}$ introduces dissipation, which is essential for breaking time-reversal symmetry and achieving nonreciprocity. Specifically, the imaginary component of $J_{3}$ encapsulates the energy exchange between the atomic ensemble and the mechanical oscillator, resulting in asymmetric optical signal transmission. From a physical standpoint, nonreciprocity arises from interference between two optical paths. One path is delineated along $a_{1}\rightarrow d\rightarrow b\rightarrow a_{2}$, while the other follows $a_{1}\rightarrow a_{2}$. The dissipative coupling $J_{3}$ enables unidirectional energy dissipation, thereby preventing back-propagation and facilitating nonreciprocal transmission.
@@ III.2 Influence of Coupling Strength and Detuning
We now examine how varying the coupling strengths $J_{2}$ and $J_{3}$, as well as the detuning $\Delta_{j}$, affect the transmission properties of the system. The study presented in Figure 3 shows the behavior of transmission amplitudes $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ as functions of the detuning $y=\delta-\omega_{m}$ for different phase difference values $\theta$ and $\phi$. It is observed that when $\theta=\phi=m\pi$, where $m\in\mathbb{Z}$, the system exhibits reciprocal behavior, characterized by equal transmission amplitudes ($\mathcal{T}_{12}=\mathcal{T}_{21}$). This phenomenon occurs because the phase differences deplete the interference effect between the two optical paths. Conversely, when $\theta\neq m\pi$, the system manifests nonreciprocity with $\mathcal{T}_{12}\neq\mathcal{T}_{21}$. Optimal nonreciprocal transmission is achieved at resonance when $\theta=\phi=\frac{\pi}{2}$, where $\mathcal{T}_{12}\approx 0$ and $\mathcal{T}_{21}\approx 1$, or when $\theta=\phi=\frac{3\pi}{2}$, where $\mathcal{T}_{12}\approx 1$ and $\mathcal{T}_{21}\approx 0$. These findings illustrate that the direction of optical signal transmission can be regulated by adjusting the phase differences between the subsystems. Specifically, for $0<\theta<\pi$ and $0<\phi<\pi$, the observation of $\mathcal{T}_{12}<\mathcal{T}_{21}$ implies stronger forward transmission from cavity 2 to cavity 1. Conversely, for $\pi<\theta<2\pi$ and $\pi<\phi<2\pi$, the reverse holds, with $\mathcal{T}_{12}>\mathcal{T}_{21}$ indicating enhanced transmission in the opposite direction.
@@
Further investigation focuses on how variations in the coupling constants $|J_{2}|$ and $J_{3}$ affect the nonreciprocal response. As demonstrated in Figure 4, increasing the coupling constant $|J_{2}|$, which regulates the interaction between the atomic ensemble and the first cavity mode, primarily affects the transmission amplitude $\mathcal{T}_{12}$. Alternatively, decreasing the coupling constant $J_{3}$ between the atomic ensemble and the mechanical mode results in the reversal of input signal transmission, as illustrated in Figure 4(g-h).
These results underscore the critical reliance of nonreciprocity on the coupling constants. When $J_{3}$ is adjusted to possess a significant imaginary component, dissipation assumes a dominant role, inhibiting backscattered signals and facilitating strong nonreciprocal transmission. Conversely, reducing the dissipation (i.e., decreasing $\Im(J_{3})$) mitigates the nonreciprocal effect, prompting the system to display more reciprocal behavior.
@@ III.3 Conditions for Perfect Nonreciprocity
We proceed to examine the required conditions for achieving perfect nonreciprocal transmission within the hybrid optomechanical system. Perfect nonreciprocity is realized when the transmission amplitude in one direction reaches unity ($\mathcal{T}_{12}=1$), while in the reverse direction it is completely suppressed ($\mathcal{T}_{21}=0$), or vice versa. This functionality is fundamental to applications in optical isolation and circulators within quantum information processing systems, where unidirectional light transmission is critical.
Perfect nonreciprocal transmission can be realized through strategic adjustments of several parameters: (i) the phase difference $\theta$ between the effective optomechanical couplings; (ii) the coupling strength $J_{3}$ between the atomic ensemble and the mechanical oscillator, with particular attention to its complex nature, where the real part signifies conservative interactions and the imaginary part indicates dissipation; (iii) the relative coupling strength $|J_{2}|$, which facilitates the interaction between the atomic ensemble and cavity 1; and (iv) the dissipation rates of the mechanical oscillator and the atomic ensemble, represented by $\gamma$ and $f$, respectively. The induction of sufficient dissipation through the imaginary part of $J_{3}$ is essential to disrupting time-reversal symmetry.
For simplicity, we consider the scenario where the phase differences are configured to $\theta=\phi=\frac{\pi}{2}$, a setup that generally yields pronounced nonreciprocal behavior. By defining $G_{1}=\sqrt{\gamma\kappa_{1}}$ and $G_{2}=\sqrt{\gamma\kappa_{2}}$, the necessary coupling strengths are determined by:
@@
The expressions above are obtained with the help of the conditions $\mathcal{T}_{12}=1$ and $\mathcal{T}_{21}=0$ (the expressions of $R_{7},R_{8},$ and $R_{9}$ are presented in Appendix). The dissipation introduced via the imaginary part of $J_{3}$ is instrumental in realizing perfect nonreciprocity. This dissipative interaction enables asymmetric energy exchange between the atomic ensemble and the mechanical oscillator, effectively mitigating back-reflections and facilitating unidirectional transmission. Conversely, reducing the dissipation (i.e., minimizing $\Im(J_{3})$) diminishes the nonreciprocal effect and tends to reinstate reciprocal transmission, where $\mathcal{T}_{12}=\mathcal{T}_{21}$.
@@ III.4 Influence of Atomic Ensemble Decay and Mechanical Damping
In order to further understand the impact of system parameters on nonreciprocal transmission, this study investigates the influence of the atomic ensemble decay rate $f$ and the mechanical damping rate $\gamma$ on the transmission amplitudes $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$. These parameters dictate the dissipation within the atomic and mechanical subsystems, respectively, and significantly affect the stability of nonreciprocity. Figures 5 and 6 illustrate that perfect nonreciprocity is maintained across a spectrum of atomic decay rates $f$, indicating that minor fluctuations in the decay rate do not substantially alter the transmission characteristics. Indeed, Figures 5(a) and 6(a) demonstrate that at reduced decay rates $f/\kappa_{2}=0.1$, the transmission amplitudes $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ show pronounced peaks in proximity to resonance. Nonetheless, nonreciprocal transmission is confined to a limited detuning range, suggesting heightened sensitivity to slight detuning variations when the atomic decay rate is minimal. With an increase in the atomic ensemble decay rate to $f/\kappa_{2}=1$, as depicted in Figures 5(b) and 6(b), the range of detuning allowing for nonreciprocal transmission expands. This results in an amplification of transmission $\mathcal{T}_{12}$ while $\mathcal{T}_{21}$ decreases markedly, indicating enhanced nonreciprocity. This expanded detuning range suggests that the system is more resilient to changes in detuning. For elevated atomic decay rates ($f/\kappa_{2}=5$), as shown in Figures 5(c) and 6(c), the system approaches near-perfect nonreciprocity over a broader detuning range. The peak magnitudes of $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ become less pronounced, implying that the system can maintain nonreciprocity. This infers that dissipation introduced by the atomic ensemble contributes a stabilizing effect in preserving nonreciprocal transmission. The findings from Figures 5 and 6 reveal that increased atomic ensemble decay rates $f$ yield wider nonreciprocal transmission bands, thereby strengthening the system’s resilience to detuning. This is consistent with the idea that increased dissipation helps suppressing back-reflections and fosters stronger nonreciprocity.
@@
Next, we investigate the influence of mechanical damping $\gamma$ on the transmission properties. Figures 7 and 8 present the transmission amplitudes $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ as functions of detuning for different values of the mechanical damping rate $\gamma$. The mechanical damping rate $\gamma$ dictates the rate at which energy is dissipated in the mechanical oscillator. Upon varying $\gamma$, we observe the interesting behaviors in the transmission amplitudes. Indeed, for low mechanical damping rates ($\gamma/\kappa_{2}=0.001$, Figures 7(a) and 8(a)), the transmission amplitudes exhibit sharp peaks around resonance. The detuning range conducive to nonreciprocal transmission is narrow, rendering the system sensitive to minor variations in detuning. In this regime, mechanical damping is insufficient to substantially influence nonreciprocity. As the mechanical damping rate increases to $\gamma/\kappa_{2}=0.01$ (Figures 7(b) and 8(b)), there is a broadening of the nonreciprocal transmission range. The peaks of $\mathcal{T}_{12}$ become broader, and $\mathcal{T}_{21}$ decreases, indicating enhanced nonreciprocity. Mechanical damping assumes a more substantial role in modulating the system’s behavior, effectively suppressing back-propagating signals. At higher mechanical damping rates ($\gamma/\kappa_{2}=0.1$ and $\gamma/\kappa_{2}=1$, Figures 7(c-d) and 8(c-d)), the detuning range over which nonreciprocity occurs is significantly broadened. The peaks of $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ become flattened, suggesting that the system may uphold nonreciprocity over an extensive detuning range. In this regime, mechanical damping exerts sufficient influence to dominate the system’s dynamics, culminating in robust nonreciprocal transmission that is less susceptible to detuning variations.
The findings from Figures 7 and 8 illustrate that an increase in mechanical damping enhances the robustness of nonreciprocity by expanding the detuning range for nonreciprocal transmission. This suggests that mechanical dissipation serves as a stabilizing function, akin to atomic ensemble decay, in sustaining nonreciprocal transmission. Overall, the above results illustrate that through modulation of phase differences, coupling strengths, and dissipation rates, it is feasible to achieve and regulate perfect nonreciprocity in the hybrid optomechanical system. These outcomes provide a foundation for the design of nonreciprocal optical devices with implications in photonics and quantum information processing.

### v2

@@ III Results and discussion
In this section, we analyze the conditions under which optical nonreciprocity can be achieved in the hybrid optomechanical system. We investigated the effects of key system parameters, such as coupling strengths, detuning, dissipation, and phase differences, on the transmission properties. The results reveal how tuning these parameters allows for control over nonreciprocal transmission, offering potential applications in optical isolation and signal routing.
@@ III.1 Optical nonreciprocal transmission
The transmission amplitudes $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ are defined to delineate the transmission of optical signal from cavity 1 to cavity 2 and vice versa. These transmission coefficients are derived from the input-output relations and exhibit a strong dependence on the system parameters, particularly the phase differences between the effective optomechanical couplings and the coupling strengths of the atomic ensemble. Figure 2 displays the dependence of transmission amplitudes $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ on the phase difference $\theta$ between the effective optomechanical couplings and the phase $\phi$ associated with the atomic ensemble. In particular, although transmission amplitudes are not markedly influenced by phase $\phi$, both phases play a crucial role in the emergence of nonreciprocal transmission.
Figure 2: Density plot of transmission amplitudes ($\mathcal{T}_{12}$ and $\mathcal{T}_{21}$) versus relative phase difference $\theta$ between effective optomechanical coupling strength and the phase $\phi$. We have used $G_{1}/\gamma=0.5$, $G_{2}=G_{1}$, $f/\gamma=10$, $\kappa_{1}/\gamma=1$, $\kappa_{2}=\kappa_{1}$, $J_{1}/\gamma=0.5$, $\omega_{m}/\gamma=10$, $\Delta^{\prime}_{1}=\Delta^{\prime}_{2}=\Delta_{\rm en}=\omega_{m}$, $|J_{2}|/\gamma=0.01$, and $J_{3}/\gamma=4.476e^{i\frac{\pi}{2}}$.
The results in Figure 2 suggest that an optimal nonreciprocal transmission is achieved when the coupling constant between the atomic ensemble and the mechanical oscillator, $J_{3}$, is complex. The complex nature of $J_{3}$ introduces dissipation, which is essential to break the time-reversal symmetry and achieve nonreciprocity. Specifically, the imaginary component of $J_{3}$ encapsulates the energy exchange between the atomic ensemble and the mechanical oscillator, resulting in asymmetric transmission of optical signals.
From a physical point of view, nonreciprocity arises from interference between two optical paths. One path is delineated along $a_{1}\rightarrow d\rightarrow b\rightarrow a_{2}$, while the other follows $a_{1}\rightarrow a_{2}$. Dissipative coupling $J_{3}$ enables unidirectional energy dissipation, thereby preventing backpropagation and facilitating nonreciprocal transmission.
@@ III.2 Influence of coupling strength and detuning
We now examine how the variation of the coupling strengths $J_{2}$ and $J_{3}$, as well as the detuning $\Delta_{j}$, affects the transmission properties of the system. The study presented in Figure 3 shows the behavior of transmission amplitudes $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ as functions of detuning $y=\delta-\omega_{m}$ for different values of phase difference $\theta$ and $\phi$. It is observed that when $\theta=\phi=m\pi$, where $m\in\mathbb{Z}$, the system exhibits reciprocal behavior, characterized by equal transmission amplitudes ($\mathcal{T}_{12}=\mathcal{T}_{21}$). This phenomenon occurs because the phase differences deplete the interference effect between the two optical paths. In contrast, when $\theta\neq m\pi$, the system manifests nonreciprocity with $\mathcal{T}_{12}\neq\mathcal{T}_{21}$. Optimal nonreciprocal transmission is achieved at resonance when $\theta=\phi=\frac{\pi}{2}$, where $\mathcal{T}_{12}\approx 0$ and $\mathcal{T}_{21}\approx 1$, or when $\theta=\phi=\frac{3\pi}{2}$, where $\mathcal{T}_{12}\approx 1$ and $\mathcal{T}_{21}\approx 0$. These findings illustrate that the direction of optical signal transmission can be regulated by adjusting the phase differences between subsystems. Specifically, for $0<\theta<\pi$ and $0<\phi<\pi$, the observation of $\mathcal{T}_{12}<\mathcal{T}_{21}$ implies stronger forward transmission from cavity 2 to cavity 1. Conversely, for $\pi<\theta<2\pi$ and $\pi<\phi<2\pi$, the reverse holds, with $\mathcal{T}_{12}>\mathcal{T}_{21}$ indicating enhanced transmission in the opposite direction.
@@
Further investigation focuses on how variations in the coupling constants $|J_{2}|$ and $J_{3}$ affect the nonreciprocal response. As demonstrated in Figure 4, increasing the coupling constant $|J_{2}|$, which regulates the interaction between the atomic ensemble and the first cavity mode, primarily affects the transmission amplitude $\mathcal{T}_{12}$. Alternatively, decreasing the coupling constant $J_{3}$ between the atomic ensemble and the mechanical mode results in a reverse transmission of the input signal, as illustrated in Figure 4 (gh).
These results underscore the critical reliance of nonreciprocity on the coupling constants. When $J_{3}$ is adjusted to possess a significant imaginary component, dissipation assumes a dominant role, inhibiting backscattered signals and facilitating strong nonreciprocal transmission. In contrast, reducing dissipation (that is, decreasing $\Im(J_{3})$) mitigates the nonreciprocal effect, prompting the system to display more reciprocal behavior.
@@ III.3 Conditions for perfect nonreciprocity
We proceed to examine the conditions required to achieve perfect nonreciprocal transmission within the hybrid optomechanical system. The perfect nonreciprocity is realized when the transmission amplitude in one direction reaches unity ($\mathcal{T}_{12}=1$), while in the reverse direction it is completely suppressed ($\mathcal{T}_{21}=0$), or vice versa. This functionality is fundamental to applications in optical isolation and circulators within quantum information processing systems, where unidirectional light transmission is critical.
Perfect nonreciprocal transmission can be realized through strategic adjustments of several parameters: (i) the phase difference $\theta$ between the effective optomechanical couplings; (ii) the coupling strength $J_{3}$ between the atomic ensemble and the mechanical oscillator, with particular attention to its complex nature, where the real part signifies conservative interactions and the imaginary part indicates dissipation; (iii) the relative coupling strength $|J_{2}|$, which facilitates the interaction between the atomic ensemble and cavity 1; and (iv) the dissipation rates of the mechanical oscillator and the atomic ensemble, represented by $\gamma$ and $f$, respectively. Induction of sufficient dissipation through the imaginary part of $J_{3}$ is essential to disrupt the symmetry of time-reversal.
For simplicity, we consider the scenario where the phase differences are configured to $\theta=\phi=\frac{\pi}{2}$, a setup that generally produces pronounced nonreciprocal behavior. By defining $G_{1}=\sqrt{\gamma\kappa_{1}}$ and $G_{2}=\sqrt{\gamma\kappa_{2}}$, the necessary coupling strengths are determined by:
@@
The expressions above are obtained with the help of conditions $\mathcal{T}_{12}=1$ and $\mathcal{T}_{21}=0$ (the expressions of $R_{7},R_{8},$ and $R_{9}$ are presented in the appendix).
The dissipation introduced through the imaginary part of $J_{3}$ is instrumental in achieving perfect nonreciprocity. This dissipative interaction enables asymmetric energy exchange between the atomic ensemble and the mechanical oscillator, effectively mitigating back-reflections and facilitating unidirectional transmission. Conversely, reducing the dissipation (i.e., minimizing $\Im(J_{3})$) diminishes the nonreciprocal effect and tends to reinstate reciprocal transmission, where $\mathcal{T}_{12}=\mathcal{T}_{21}$.
@@ III.4 Influence of atomic ensemble decay and mechanical damping
To better understand the impact of system parameters on nonreciprocal transmission, this study investigates the influence of the decay rate of the atomic ensemble $f$ and the mechanical damping rate $\gamma$ on transmission amplitudes $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$. These parameters dictate the dissipation within the atomic and mechanical subsystems and significantly affect the stability of nonreciprocity. Figures 5 and 6 illustrate that perfect nonreciprocity is maintained throughout a spectrum of atomic decay rates $f$, indicating that minor fluctuations in the decay rate do not substantially alter the transmission characteristics. In fact, Figures 5 (a) and 6 (a) demonstrate that at reduced decay rates $f/\kappa_{2}=0.1$, the transmission amplitudes $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ show pronounced peaks near the resonance. However, nonreciprocal transmission is confined to a limited detuning range, suggesting heightened sensitivity to slight detuning variations when the atomic decay rate is minimal. With an increase in the decay rate of the atomic ensemble to $f/\kappa_{2}=1$, as depicted in Figures 5(b) and 6(b), the range of detuning allowing nonreciprocal transmission expands. This results in amplification of transmission $\mathcal{T}_{12}$ while $\mathcal{T}_{21}$ decreases markedly, indicating enhanced nonreciprocity. This expanded detuning range suggests that the system is more resilient to changes in detuning. For elevated atomic decay rates ($f/\kappa_{2}=5$), as shown in Figures 5(c) and 6(c), the system approaches near-perfect nonreciprocity over a broader detuning range. The peak magnitudes of $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ become less pronounced, implying that the system can maintain nonreciprocity. This infers that dissipation introduced by the atomic ensemble contributes to a stabilizing effect in preserving nonreciprocal transmission. The findings of Figures 5 and 6 reveal that the higher decay rates of the atomic ensemble $f$ produce wider nonreciprocal transmission bands, thus strengthening the system’s resilience to detuning. This is consistent with the idea that increased dissipation helps suppressing back-reflections and fosters stronger nonreciprocity.
@@
Next, we investigate the influence of mechanical damping $\gamma$ on transmission properties. Figures 7 and 8 present transmission amplitudes $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ as functions of detuning for different values of the mechanical damping rate $\gamma$. The mechanical damping rate $\gamma$ dictates the rate at which energy is dissipated in the mechanical oscillator. Upon varying $\gamma$, we observe the interesting behaviors in the transmission amplitudes. In fact, for low mechanical damping rates ($\gamma/\kappa_{2}=0.001$, Figures 7(a) and 8(a)), the transmission amplitudes exhibit sharp peaks around the resonance. The detuning range conducive to nonreciprocal transmission is narrow, making the system sensitive to minor variations in detuning. In this regime, mechanical damping is insufficient to substantially influence nonreciprocity. As the mechanical damping rate increases to $\gamma/\kappa_{2}=0.01$ (Figures 7(b) and 8(b)), there is a broadening of the nonreciprocal transmission range. The peaks of $\mathcal{T}_{12}$ become wider and $\mathcal{T}_{21}$ decreases, indicating enhanced nonreciprocity. Mechanical damping assumes a more substantial role in modulating the system’s behavior, effectively suppressing back-propagating signals. At higher mechanical damping rates ($\gamma/\kappa_{2}=0.1$ and $\gamma/\kappa_{2}=1$, Figures 7(c-d) and 8(c-d)), the detuning range over which nonreciprocity occurs is significantly broadened. The peaks of $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ flatten, suggesting that the system may maintain nonreciprocity over an extensive detuning range. In this regime, mechanical damping exerts sufficient influence to dominate the system’s dynamics, culminating in robust nonreciprocal transmission that is less susceptible to detuning variations.
The findings of Figures 7 and 8 illustrate that an increase in mechanical damping enhances the robustness of nonreciprocity by expanding the de-tuning range for nonreciprocal transmission. This suggests that mechanical dissipation serves as a stabilizing function, akin to the decay of the atomic ensemble, in the maintenance of nonreciprocal transmission. Overall, the results presented above illustrate that, through modulation of phase differences, coupling strengths, and dissipation rates, it is feasible to achieve and regulate perfect nonreciprocity in the hybrid optomechanical system.
Figure 9: Transmission amplitudes $\mathcal{T}_{12}$ (black solid–line) and $\mathcal{T}_{21}$ (blue dotted–line) as a function of detuning $y/\gamma$ for different values of $J_{3}$. Parameters used are the same as for Figure 2 except for $\theta=\phi=\frac{\pi}{2}$ and complex coupling $J_{3}$ (a) $J_{3}/\gamma=4.676e^{i\frac{\pi}{2}}$, (b) $J_{3}/\gamma=4.276e^{i\frac{\pi}{2}}$
These outcomes provide a foundation for the design of nonreciprocal optical devices with implications in photonics and quantum information processing.
It can be clearly seen from Figure 9, that our system can show gain (amplification) and loss (decay) by carefully adjusting the value of the complex coupling constant $J_{3}$.